WAC Notes August 27 2019

From PREX Wiki
Jump to: navigation, search
Documentation
HOW TOs for shift crew
Expert Tools
All Expert Contacts



PREX Main << Weekly Analysis Coordinator << WAC Notes << Instructions

WAC Notes August 26 2019 << >> WAC Notes August 28 2019

August 27 2019

Daily Meeting BlueJeans

Production Run List - Slug 70

  • Run 4585
    • BCM double peaking on double difference
  • Run 4586
    • Large difference between BPM4A-X and its complement
  • Run 4587
  • Run 4588
  • Run 4589
    • BCM double differences is still weirdly peaked, not terribly so
  • Run 4590
  • Run 4591
  • Run 4592
  • Run 4593
  • Run 4594
  • Run 4595

Production Run List - Slug 71

So this slug had a lot of problems with widths due in part to a unchecked raster size and unsuitable optics.

CONFIG CHANGE: FFB On, Aq ON, BMW On, FF On

  • Run 4596
    • Small BPM drifts on 4aX, advise doing nothing

CONFIG CHANGE: FFB Off, Aq On, BMW On, FF On

  • Run 4621
    • Main Detectors
      • Widths finally down again
    • BCMs
      • BCM wings again
  • Run 4622
  • BCMs
    • Wings

CONFIG CHANGE: FFB Off, Aq On, BMW On

Non-Production Runs and Bad Runs

The following runs had bad widths due to the raster and optics being bad

  • Run 4606
    • Main Dets
      • Widths untenable
    • BPMs
      • BPM drifts at start of run, but in all honesty, does it matter?
  • Run 4607
    • Main Dets
      • Widths bad
  • Run 4608
    • Four seconds long
  • Run 4609
    • Main Dets
      • Widths Bad
      • BPM drifts at start and end of runs

CONFIG: FFB Off, Aq On, BMW On, FF Off

  • Run 4612
    • Main Dets
      • Widths bad
    • BPMs
      • BPM 4aY drift at end of run >1 mm
      • BPM 4ec IDFK
  • Run 4613
    • Main Dets
      • Widths bad
    • BPMs
      • BPM 4aY, 4eY drifts at end of run
      • 4ec weird behavior

CONFIG: FFB On, Aq On, BMW On

  • Run 4614
    • Main Dets
      • Widths bad
    • BPMs
      • 4a, 4e drifts at start of run
  • Run 4615
    • Main Dets
      • Widths bad
  • Run 4616
    • Main Dets
      • Widths bad
  • Run 4617
    • Main Dets
      • Widths still kinda bad?
    • BPMs
      • BPM drifts in both 4a, 4e, and weird behavior in 4ac, and 4ec
  • Run 4618
    • Main Dets
      • Widths bad
    • BCMs
      • Long beam down period in middle of run
    • BPMs
      • Beam drifts
    • Compton re-tuning happened during run
  • Run 4619
    • Missing?
  • Run 4620
    • Main Dets
      • Widths still high
    • BPMs
      • BPM drifts, not that bad


Pedestal/Calibrations Updates

  • Run 4610: Calibration
  • Run 4611: PITA Scan

Issues

Other Status Updates